When the price of movies decreases from $15 to $10, the number of tickets sold increases from 3000 to 5000. What is the price elasticity of demand using the midpoint formula? Select one O a. 400 O b. 1.8 c. 1.25 O d.0.025
Answer #1

P1 =$15

P2 =$10

Q1 =3000

Q2 =5000

By mid-point formula:

Price elasticity of demand = (% change in quantity demanded )/ (% change in price)

% change in quantity demanded = (5000-3000)/ [(5000+3000)/2] 100 = (2000/ 4000) 100=(0.5)100= 50%

% change in price = (10-15)/ [(10+15)/2] 100 = (-5/12.5) 100= -0.4(100) = -40%

Price elasticity of demand = (50/-40)= -1.25

Hence, option(C) is correct.